ratio of the length to width of a rectangle is 3:2.
The perimeter is 70 cm.
Whar are the dimensions of the rectangle?

Answer by FireFox(3)

w = width l = length
3l + 2w = 70
3l = 2w
---------------
2w + 2w = 70
=4w = 70
=w = 17.5
{Width = 17.5}
--------------
3l + 2 x 17.5 = 70
31+ 35 = 70
3l = 70-35
3l = 35
l = 11.666

w= 17.5 l = 11.67